MICROWAVE SPECTRA OF THE METHYLCYANOPOLYYNES CH_{3}(C ≡C)_{n}CN, n=2,3,4,5

WEI CHEN, JENS-UWE GRABOW, MICHAEL J. TRAVERS, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138; and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138; \underline{STEWART E. NOVICK}, MICHAELEEN R. MUNROW, Department of Chemistry, Wesleyan University, Middletown, CT 06459; M. C. McCARTHY, P. THAD-DEUS, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138; and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138.

The ground state rotational spectra of four methylcyanopolyynes, CH_{3}(C ≡C)_{2}CN, CH_{3}(C ≡C)_{3}CN, CH_{3}(C ≡C)_{4}CN, and CH_{3}(C ≡C)_{5}CN, were measured between 6 and 22 GHz by Fourier transform microwave spectroscopy. For CH_{3}(C ≡C)_{2}CN, 8 transitions, each with nitrogen nuclear quadrupole hyperfine structure (hfs), were observed and a value of eqQ = – 4.25(3) MHz was determined for the first time. Twenty-one rotational transitions of CH_{3}(C ≡C)_{3}CN, ten rotational transitions of CH_{3}(C ≡C)_{4}CN, and eleven rotational transitions of CH_{3}(C≡C)_{5}CN were detected for the first time and precise values for the rotational, leading centrifugal distortion constants of each molecule, and the nuclear quadrupole coupling constant of CH_{3}(C ≡C)_{3}CN were determined.